# Towards Machine-Generated Code for the Resolution of User Intentions

### v1

@@ Abstract
The growing capabilities of Artificial Intelligence (AI), particularly Large Language Models (LLMs), prompt a reassessment of the interaction mechanisms between users and their devices. Currently, users are required to use a set of high-level applications to achieve their desired results. However, the advent of AI may signal a shift in this regard, as its capabilities have generated novel prospects for user-provided intent resolution through the deployment of model-generated code, which is tantamount to the generation of workflows comprising a multitude of interdependent steps. This development represents a significant progression in the realm of hybrid workflows, where human and artificial intelligence collaborate to address user intentions, with the former responsible for defining these intentions and the latter for implementing the solutions to address them. In this paper, we investigate the feasibility of generating and executing workflows through code generation that results from prompting an LLM with a concrete user intention, such as Please send my car title to my insurance company, and a simplified application programming interface for a GUI-less operating system. We provide in-depth analysis and comparison of various user intentions, the resulting code, and its execution. The findings demonstrate a general feasibility of our approach and that the employed LLM, GPT-4o-mini, exhibits remarkable proficiency in the generation of code-oriented workflows in accordance with provided user intentions.
@@ 3 Preliminary Considerations
From a macro perspective, a single workflow can be conceptualized as a finite state machine, constituted of multiple steps, which must be executed in a predetermined and ordered fashion for the resolution of a user intention. The interconnection of these steps allows for control flow with loops as well as case differentiations. Ideally, individual steps of these workflows are reusable throughout different workflows for reducing the overall complexity. Figure 1 illustrates an exemplary schematic of such a workflow generated from an intention by an AI from a semantic point of view.
Figure 1: Exemplary schematic showing a workflow with multiple asynchronous and eventually recurring steps.
The elements of workflows and their depiction as finite state machines, in conjunction with the employment of control flow with loops and case differentiations, directly result in the consideration and utilisation of programming languages in the context of workflows and their execution. LLMs can support this execution model by their ability to generate code.
@@ 4 Method
In order to elucidate the concrete relationship between the components presented in our system architecture, we present an exemplar data flow from intention to the execution of the generated workflow.

### v2

@@ Abstract
The growing capabilities of Artificial Intelligence (AI), particularly Large Language Models (LLMs), prompt a reassessment of the interaction mechanisms between users and their devices. Currently, users are required to use a set of high-level applications to achieve their desired results. However, the advent of AI may signal a shift in this regard, as its capabilities have generated novel prospects for user-provided intent resolution through the deployment of model-generated code. This development represents a significant progression in the realm of hybrid workflows, where human and artificial intelligence collaborate to address user intentions, with the former responsible for defining these intentions and the latter for implementing the solutions to address them. In this paper, we investigate the feasibility of generating and executing workflows through code generation that results from prompting an LLM with a concrete user intention, and a simplified application programming interface for a GUI-less operating system. We provide an in-depth analysis and comparison of various user intentions, the resulting code, and its execution. The findings demonstrate the general feasibility of our approach and that the employed LLM, GPT-4o-mini, exhibits remarkable proficiency in the generation of code-oriented workflows in accordance with provided user intentions.
@@ 3 Preliminary Considerations
From a macro perspective, a single workflow can be conceptualized as a finite state machine, constituted of multiple steps, which must be executed in a predetermined and ordered fashion for the resolution of a user intention. The interconnection of these steps allows for control flow with loops as well as case differentiations. Ideally, individual steps of these workflows are reusable throughout different workflows for reducing the overall complexity. The elements of workflows and their depiction as finite state machines, in conjunction with the employment of control flow with loops and case differentiations, directly result in the consideration and utilisation of programming languages in the context of workflows and their execution. LLMs can support this execution model by their ability to generate code.